I am wondering if anyone could help me with this story problem. Here are the directions: Write and solve an equation, inequality, or system of equations to answer the question.
Here is the problem:
Olivia bought 5 greeting cards that were each the same price. She used coupons worth $2.50. Alexandra bought 4 cards and used a $1.00 coupon. Each girl paid the same amount. What was the price of each card?


Answer by mananth(16946)

Here is the problem:
Olivia bought 5 greeting cards that were each the same price. She used coupons worth $2.50. Alexandra bought 4 cards and used a $1.00 coupon. Each girl paid the same amount. What was the price of each card?
Let the price of the card be $x
Olivia
5 cards = 5x ---coupons worth 2.50
So she paid 5x-2.5
Alexandria
4 cards so cost = 4x
coupons =$1
4x-1
they both paid same amount
5x-2.5=4x-1
5x-4x=2.5-1
x=1.5
Card costs $ 1.50
Ananth